# The Open Journal of Astrophysics

### v1

@@ 1 Introduction
In the universe, galaxy clusters are the largest structures bound by their own gravity, centered at the maxima of the universal density field. With large reservoirs of hot gas and deep gravitational potentials, the cluster environment significantly shapes the evolution and formation of nearby galaxies. Hence, galaxy clusters serve as excellent cosmic laboratories for the study of galaxy formation and evolutions. To adequately use observations to study galaxy formation, understanding the spatial extent of a cluster is crucial, as it allows observers to conveniently separate cluster and field populations based on their positions relative to the closest cluster environment. Such transition is delineated by the evolution of intrinsic galaxy properties within the cluster environment a galaxy resides in.
@@
Observations reveal that the cluster and field environments impact galaxy formation in different ways. Particularly, cluster galaxies are older, redder, and more elliptical, while their field counterparts tend to be bluer and more spiral (Dressler, 1980; Cooper et al., 2006; Donnari et al., 2021). Therefore, residing in different environments can significantly alter the trajectory of a galaxy’s evolution. Previous studies have investigated the temporal impact of the cluster environment, as well as the spatial extent of mergers and mass accretion in galaxy cluster outskirts (e.g., Cuesta et al., 2008; Behroozi et al., 2014; O’Neil et al., 2024). However, a systematic study has not been conducted to examine the spatial transition marked by the way intrinsic galaxy properties respond to the cluster and field environment. As a galaxy enters the cluster environment, it experiences a series of interactions with intracluster medium of the host cluster (ICM). The truncation of a galaxy’s star formation, for example, is often referred to as "quenching", during which numerous processes occur simultaneously, making it difficult to unravel. Quenching can occur both internally as mass quenching and globally as environmental quenching. The former often refers to processes within the galaxies, such as Active Galactic Nuclei feedback (Croton et al., 2006) and Supernovae (Larson, 1974). One type of environmental quenching is ram pressure stripping, where the gas of a galaxy is stripped away by the intracluster medium, making its star formation less active compared to galaxies in the field (Gunn & Gott, 1972; Abadi, Moore & Bower, 1999). This process is governed by the galaxy’s orbital parameters and intrinsic properties, as well as the physical conditions of the host cluster’s environment. Other intracluster activities such as strangulation (Tinsley & Caldwell, 1980), the high‐speed galaxy–galaxy “harassment” (Moore et al., 1996), and tidal disruption (Merritt, 1983) all contribute to the quenching of star formation in galaxies. According to (Woo & Cooper, 2012), the quenching of galaxies by the halo-environment at $z=0$ mostly depends on dark matter halo mass. The quenched fraction is reported to increase with halo mass and stellar mass. Moreover, cosmological simulations suggest that hot gas reservoirs in halos of mass $\gtrsim 10^{12}\,{\rm M}_{\odot}$ drive both mass and environment quenching (Gabor & Davé, 2015). Observations of clusters up to $z\approx 1$ show that quenching efficiency depends jointly on galaxy stellar mass and environment, with both effects being inseparable in the densest regions (Pintos-Castro et al., 2019). Semi-analytic modeling further reveals that stellar mass is the primary quenching mechanism, while environmental processes modulate the quenched fraction via halo mass and local density (Contini et al., 2019). These elements involved in this process give rise to the opportunity to classify galaxies based on their cluster memberships, where galaxies that have not undergone such impacts may exhibit a different set of patterns in their intrinsic properties. As a result, the spatial extent of the cluster environment’s influence can be informed by galaxy memberships.
@@
Observations and simulations reveal that baryon fractions within these clusters tend to stabilise early on, especially in massive ones (Chiu et al., 2018; Angelinelli et al., 2023). This suggests that the co-evolution of galaxies and dark matter haloes can by inferred from the environmental conditions encountered by cluster members, with minimal complication from evolving baryonic contents. This provides an opportunity to investigate how galaxies transition near cluster outskirts, particularly the spatial distribution of properties that reflect environmental influence. One key insight is the shift in the probability of a galaxy being classified as a field or cluster member with distance, which can be informed by measurable characteristics. Baryonic properties, such as gas fraction, colour, stellar-to-subhalo mass ratio, metallicity and star formation rate, as well as dark matter-related quantities like subhalo dark matter mass and local density, serve as valuable metrics for identifying environmental transitions and studying the transition region between field and cluster galaxies (e.g., Wetzel et al., 2012; Haines et al., 2015; Oxland et al., 2024). While it is common to infer galactic properties based on observable signatures, systematic studies that explain how the cluster environment affects galaxy membership remain limited. Such studies should ideally delineate the transition region between field galaxies and cluster galaxies, as these two categories exhibit markedly different intrinsic properties. Large-volume cosmological simulations of galaxy formation offer the opportunity to explore these transitions in detail, as they reproduce a wide range of galaxy characteristics and scaling relations representative of real-world observations.
@@
Previously studies have established cluster boundary definitions that inform our work. A physically motivated boundary that marks the onset of the cluster environment’s dynamical structure and influence is the splashback radius ($R_{\mathrm{sp}}$). It defines the outer boundary of a halo separating infalling matter from collapsed, virialised material (Diemer & Kravtsov, 2014; Adhikari, Dalal & Chamberlain, 2014; More, Diemer & Kravtsov, 2015). This radius corresponds to a sharp steepening in the halo density profile, occurring at the location where the logarithmic slope reaches a minimum. The splashback surface marks the transition between single-stream infall and multi-stream, virialised flows, thus providing a dynamical boundary for the halo (Bertschinger, 1985; Vogelsberger et al., 2009), which leads them to continue evolving under the influence of the cluster environment (O’Neil et al., 2024). As the collision-less dark matter particles collapse gravitationally to form the dark matter haloes, they orbit the potential of the halo to form the multi-streaming region. Galaxies that enter the cluster environment have dynamics that follow the trajectory of the dark matter particles, which leads them to continue evolving under the influence of the cluster. This phenomenon shows that there exists a transition in the cluster environment that represents where galaxies enter the intracluster medium. This boundary distinguishes galaxy dynamics significantly. Therefore, dynamical galactic properties exhibit transitions across the splashback boundary. This is demonstrated by the first-infall trajectories of galaxies outside $R_{sp}$ being predominantly radial, whereas those inside mainly have more tangential and virialised orbits. In our study, several dynamical properties are selected to test the extent of such effect.
@@
Environmental processes within the cluster environment not only impact dynamical behaviour but also significantly influence the stellar content of galaxies. As noted by Ahad et al. (2021), environmental quenching alters the stellar mass function due to the dense surroundings and increased interactions among galaxies. Since any mechanism that perturbs or removes gas plays a pivotal role in regulating star formation, gas-related properties—i.e. gas-to-baryon mass fraction and gas metallicity—are excellent indicators of galaxy evolution (Pérez-Millán et al., 2023). Thus, beyond dynamical considerations, the cluster environment’s influence on baryonic properties must also be considered to fully understand the transition of galaxies near cluster outskirts. This emphasis is motivated by the fact that galaxy clusters, embedded in the Cosmic Web, act as large multi-phase gas reservoirs. The effect of the collisional gas component and gas accretion within the environment cannot be inferred purely from the dark matter density field (O’Kane et al., 2024). As galaxies fall into cluster cores, they encounter and interact with the hot intracluster medium, which drives significant transformations in their baryonic content. Consequently, the cluster boundary can also be traced via the behaviour of gas, offering a complementary probe. In particular, accreting gas at the cluster outskirts undergoes shock heating during its first infall. This process, marked by a sudden jump in the gas entropy profile, defines what is known as the cosmic accretion shock. It arises as low-density, pristine gas collapses from void regions onto the cluster potential, forming a physical boundary delineated by the transition to hot, collapsed gas (Lau et al., 2015). Recent studies such as (Aung, Nagai & Lau, 2021) demonstrate that the location of the accretion shock radius—identified through entropy gradients—is systematically offset from splashback-based boundaries. Specifically, the accretion shock radius tends to lie beyond the splashback radius ($R_{\rm sp}$), as well as other common definitions such as $R_{200,\mathrm{mean}}$. These findings mark the complexity of defining a universal cluster boundary. The offsets among $R_{\rm sp}$, $R_{\rm shock}$, $R_{200,\mathrm{mean}}$ and $R_{200,\mathrm{crit}}$ suggest that each boundary metric traces distinct physical processes. Therefore, one of the goals of this study is to evaluate whether these boundary definitions reliably capture the transition between cluster and field galaxies in both dark matter and baryonic contexts. Thus, due to the multitude of environmental components and the complex nature of galaxy evolution, the boundary at which galaxies transition from the field to the cluster environment is inherently non-trivial. Given such complexity of physical processes at cluster outskirts and the multi-phase structure of the cluster environment, we propose a machine learning (ML) framework to find a probabilistic transition region between cluster and field populations around cluster outskirts. The distinct physical properties and evolutionary histories of field and cluster galaxies create an opportunity for pattern recognition through ML methods. In recent years, ML techniques have seen increasing application in data-driven cosmology and astrophysics, particularly in galaxy classification tasks (e.g. Farid et al., 2023). We employ a supervised learning approach using Deep Neural Networks (DNNs) to infer the spatial extent of the cluster environment’s impact on galaxy evolution. Galaxy memberships—i.e., whether a galaxy is influenced by the cluster environment—serve as classification labels. Selected galactic intrinsic properties that are sensitive to environmental effects are used as input features for the neural network. The selection of input features plays a pivotal role in capturing the probabilistic transition across the cluster boundary. Our approach is motivated by two recent developments. First, the cluster outskirts host a kinematically mixed population of pristine infalling galaxies and processed orbiting galaxies (Aung et al., 2020). Second, the gas accretion shock, where processes like ram-pressure stripping can initiate quenching (e.g. Zinger et al., 2018), often extends significantly beyond the dark matter splashback radius (Aung, Nagai & Lau, 2021). This results in a physically consequential region where the environment is already shaping galaxy properties. We therefore need a new tool that expands beyond a hard boundary and is instead sensitive to the physical properties of galaxies themselves. Our work introduces a probabilistic boundary designed precisely to mark this transitional zone using intrinsic galaxy properties.
@@ 2.2 Galaxy and cluster samples
The TNG simulation suite identifies groups using the Friends-of-Friends (FoF) algorithm (Davis et al., 1985), where particles are grouped together by their spatial distribution. The SubFind Algorithm (Springel et al., 2001; Dolag et al., 2009) identifies subhaloes based on gravitationally bound particles. The Friends-of-Friends (FoF) and SubFind algorithms are ran by the TNG Collaboration, which produce the group catalogue used in this work. From this catalogue, we first select samples out of all the groups and subhaloes at redshift $z=0$. The selection criteria is described as follows. Haloes with $M_{200,mean}$ > $10^{13}\,{\rm M}_{\odot}$ are considered clusters, and subhaloes with $M_{\star}$ > $10^{9}\,{\rm M}_{\odot}$ and $M_{dm}$ > $10^{9.5}\,{\rm M}_{\odot}$ are considered galaxies. The thresholds for the galaxy definitions was motivated by the resolution limit, using subhalo mass functions, as described in (Nagai & Kravtsov, 2005). We then group the cluster samples into four mass bins based on their $M_{200\rm,mean}$: $10^{13\text{--}13.5}{\rm M}_{\odot}$, $10^{13.5\text{--}14}{\rm M}_{\odot}$, $10^{14\text{--}14.5}{\rm M}_{\odot}$, and $10^{14.5\text{--}15}{\rm M}_{\odot}$. These mass ranges are used for further analysis throughout the paper. They are chosen due to the lack of haloes with mass greater than $10^{15}M_{\odot}$ to form a mass bin of $10^{15-15.5}M_{\odot}$. The mass bins were chosen to contain enough cluster samples for stacking.
Even though the simulation associates all subhaloes with host groups, these host groups may not suffice to be a cluster environment. Hence, we use a sorting algorithm to match each galaxy to the nearest cluster. The algorithm is applied to all subhaloes and matches them to our selected clusters. This way, we can ensure that the selected galaxies are paired with their true closest cluster within the box, instead of only relative to the set of chosen galaxies. We calculate the point-to-point distance in a periodic box. At the end of Algorithm 1 in Appendix A, we have calculated the distance between each subhalo and its closest cluster, while logging the index of its home cluster. During implementation steps in later sections, we normalise the distance by the $R_{200,\rm mean}$ of the host cluster.

### v2

@@ 1 Introduction
In the universe, galaxy clusters are the largest structures bound by their own gravity, centered at the maxima of the universal density field. With large reservoirs of hot gas and deep gravitational potentials, the cluster environment significantly shapes the evolution and formation of nearby galaxies. Hence, galaxy clusters serve as excellent cosmic laboratories for the study of galaxy formation and evolution. To adequately use observations to study galaxy formation, understanding the spatial extent of a cluster is crucial, as it allows observers to conveniently separate cluster and field populations based on their positions relative to the closest cluster environment. Such transition is delineated by the evolution of intrinsic galaxy properties within the cluster environment a galaxy resides in.
@@
Observations reveal that the cluster and field environments impact galaxy formation in different ways. Particularly, cluster galaxies are older, redder, and more elliptical, while their field counterparts tend to be bluer and more spiral (Dressler, 1980; Cooper et al., 2006; Donnari et al., 2021). Therefore, residing in different environments can significantly alter the trajectory of a galaxy’s evolution. Previous studies have investigated the temporal impact of the cluster environment, as well as the spatial extent of mergers and mass accretion in galaxy cluster outskirts (e.g., Cuesta et al., 2008; Behroozi et al., 2014; O’Neil et al., 2024). However, a systematic study has not been conducted to examine the spatial transition marked by the way intrinsic galaxy properties respond to the cluster and field environment. As a galaxy enters the cluster environment, it experiences a series of interactions with the intracluster medium of the host cluster (ICM). The truncation of a galaxy’s star formation, for example, is often referred to as "quenching", during which numerous processes occur simultaneously, making it difficult to unravel. Quenching can occur both internally as mass quenching and globally as environmental quenching. The former often refers to processes within the galaxies, such as Active Galactic Nuclei feedback (Croton et al., 2006) and Supernovae (Larson, 1974). One type of environmental quenching is ram pressure stripping, where the gas of a galaxy is stripped away by the intracluster medium, making its star formation less active compared to galaxies in the field (Gunn & Gott, 1972; Abadi, Moore & Bower, 1999). This process is governed by the galaxy’s orbital parameters and intrinsic properties, as well as the physical conditions of the host cluster’s environment. Other intracluster activities such as strangulation (Tinsley & Caldwell, 1980), the high‐speed galaxy–galaxy “harassment” (Moore et al., 1996), and tidal disruption (Merritt, 1983) all contribute to the quenching of star formation in galaxies. According to Woo & Cooper (2012), the quenching of galaxies by the halo-environment at $z=0$ mostly depends on dark matter halo mass. The quenched fraction is reported to increase with halo mass and stellar mass. Moreover, cosmological simulations suggest that hot gas reservoirs in halos of mass $\gtrsim 10^{12}\,{\rm M}_{\odot}$ drive both mass and environment quenching (Gabor & Davé, 2015). Observations of clusters up to $z\approx 1$ show that quenching efficiency depends jointly on galaxy stellar mass and environment, with both effects being inseparable in the densest regions (Pintos-Castro et al., 2019). Semi-analytic modeling further reveals that stellar mass is the primary quenching mechanism, while environmental processes modulate the quenched fraction via halo mass and local density (Contini et al., 2019). These elements involved in this process give rise to the opportunity to classify galaxies based on their cluster memberships, where galaxies that have not undergone such impacts may exhibit a different set of patterns in their intrinsic properties. As a result, the spatial extent of the cluster environment’s influence can be informed by galaxy memberships.
@@
Observations and simulations reveal that baryon fractions within these clusters tend to stabilise early on, especially in massive ones (Chiu et al., 2018; Angelinelli et al., 2023). In observations, the baryonic fraction is measure by inferring the mass of hot gas in the intracluster medium (ICM) using observed X-ray imaging and the thermal SZ signal (Birkinshaw, 1999; LaRoque et al., 2006; Allen, Evrard & Mantz, 2011). The total mass is typically estimated using strong or weak gravitational lensing, which probes the projected mass distribution and benchmarks the calibrated cluster total mass. In simulations, all baryonic components can be directly obtained by summing up the gas and stellar contents within the cluster’s associated particles (Kravtsov & Borgani, 2012). It is important to distinguish cluster-scale baryonic fractions and galaxy-scale baryonic properties (e.g. gas fraction, SFR, and gas/stellar metallicities), where the latter is generally easier to measure in simulations (Planelles et al., 2013). The co-evolution of galaxies and dark matter haloes can be inferred from the environmental conditions encountered by cluster members, with minimal complication from evolving baryonic contents. This provides an opportunity to investigate how galaxies transition near cluster outskirts, particularly the spatial distribution of properties that reflect environmental influence. One key insight is the shift in the probability of a galaxy being classified as a field or cluster member with distance, which can be informed by measurable characteristics (Rykoff et al., 2014; Castignani & Benoist, 2016). Baryonic properties, such as gas fraction, colour, stellar-to-subhalo mass ratio, metallicity and star formation rate, as well as dark matter-related quantities like subhalo dark matter mass and local density, serve as valuable metrics for identifying environmental transitions and studying the transition region between field and cluster galaxies (e.g., Wetzel et al., 2012; Haines et al., 2015; Oxland et al., 2024). While it is common to infer galactic properties based on observable signatures, systematic studies that explain how the cluster environment affects galaxy membership remain limited. Such studies should ideally delineate the transition region between field galaxies and cluster galaxies, as these two categories exhibit markedly different intrinsic properties. Large-volume cosmological simulations of galaxy formation offer the opportunity to explore these transitions in detail, as they reproduce a wide range of galaxy characteristics and scaling relations representative of real-world observations.
@@
Previously studies have established cluster boundary definitions that inform our work. A physically motivated boundary that marks the onset of the cluster environment’s dynamical structure and influence is the splashback radius ($R_{\mathrm{sp}}$). It defines the outer boundary of a halo separating infalling matter from collapsed, virialised material (Diemer & Kravtsov, 2014; Adhikari, Dalal & Chamberlain, 2014; More, Diemer & Kravtsov, 2015). This radius corresponds to a sharp steepening in the halo density profile, occurring at the location where the logarithmic slope reaches a minimum. The splashback surface marks the transition between single-stream infall and multi-stream, virialised flows, thus providing a dynamical boundary for the halo (Bertschinger, 1985; Vogelsberger et al., 2009), which leads them to continue evolving under the influence of the cluster environment (O’Neil et al., 2024). As the collision-less dark matter particles collapse gravitationally to form the dark matter haloes, they orbit the potential of the halo to form the multi-streaming region. Galaxies that enter the cluster environment have dynamics that follow the trajectory of the dark matter particles, which leads them to continue evolving under the influence of the cluster. This phenomenon shows that there exists a transition in the cluster environment that represents where galaxies enter the intracluster medium. This boundary distinguishes galaxy dynamics significantly. Therefore, dynamical galactic properties exhibit transitions across the splashback boundary. This is demonstrated by the first-infall trajectories of galaxies outside $R_{sp}$ being predominantly radial, whereas those inside mainly have more tangential and virialised orbits (Pizzardo et al., 2024). In our study, several dynamical properties are selected to test the extent of such effect.
@@
Environmental processes within the cluster environment not only impact dynamical behaviour but also significantly influence the stellar content of galaxies. As noted by Ahad et al. (2021), environmental quenching alters the stellar mass function due to the dense surroundings and increased interactions among galaxies. Since any mechanism that perturbs or removes gas plays a pivotal role in regulating star formation, gas-related properties—i.e. gas-to-baryon mass fraction and gas metallicity—are excellent indicators of galaxy evolution (Pérez-Millán et al., 2023). Thus, beyond dynamical considerations, the cluster environment’s influence on baryonic properties must also be considered to fully understand the transition of galaxies near cluster outskirts. This emphasis is motivated by the fact that galaxy clusters, embedded in the Cosmic Web, act as large multi-phase gas reservoirs. The effect of the collisional gas component and gas accretion within the environment cannot be inferred purely from the dark matter density field (O’Kane et al., 2024). As galaxies fall into cluster cores, they encounter and interact with the hot intracluster medium, which drives significant transformations in their baryonic content. Consequently, the cluster boundary can also be traced via the behaviour of gas, offering a complementary probe. In particular, accreting gas at the cluster outskirts undergoes shock heating during its first infall. This process, marked by a sudden jump in the gas entropy profile, defines what is known as the cosmic accretion shock. It arises as low-density, pristine gas collapses from void regions onto the cluster potential, forming a physical boundary delineated by the transition to hot, collapsed gas (Lau et al., 2015). Recent studies such as Aung, Nagai & Lau (2021) demonstrate that the location of the accretion shock radius—identified through entropy gradients—is systematically offset from splashback-based boundaries. Specifically, the accretion shock radius tends to lie beyond the splashback radius ($R_{\rm sp}$), as well as other common definitions such as $R_{200,\mathrm{mean}}$. These findings mark the complexity of defining a universal cluster boundary. The offsets among $R_{\rm sp}$, $R_{\rm shock}$, $R_{200,\mathrm{mean}}$ and $R_{200,\mathrm{crit}}$ suggest that each boundary metric traces distinct physical processes. Therefore, one of the goals of this study is to evaluate whether these boundary definitions reliably capture the transition between cluster and field galaxies in both dark matter and baryonic contexts. Thus, due to the multitude of environmental components and the complex nature of galaxy evolution, the boundary at which galaxies transition from the field to the cluster environment is inherently non-trivial. Given such complexity of physical processes at cluster outskirts and the multi-phase structure of the cluster environment, we propose a machine learning (ML) framework to find a probabilistic transition region between cluster and field populations around cluster outskirts. The distinct physical properties and evolutionary histories of field and cluster galaxies create an opportunity for pattern recognition through ML methods. In recent years, ML techniques have seen increasing application in data-driven cosmology and astrophysics, particularly in galaxy classification tasks (e.g. Farid et al., 2023). We employ a supervised learning approach using Deep Neural Networks (DNNs) to infer the spatial extent of the cluster environment’s impact on galaxy evolution. Galaxy memberships—i.e., whether a galaxy is influenced by the cluster environment—serve as classification labels. Selected galactic intrinsic properties that are sensitive to environmental effects are used as input features for the neural network. The selection of input features plays a pivotal role in capturing the probabilistic transition across the cluster boundary. Our approach is motivated by two recent developments. First, the cluster outskirts host a kinematically mixed population of pristine infalling galaxies and processed orbiting galaxies (Aung et al., 2020). Second, the gas accretion shock, where processes like ram-pressure stripping can initiate quenching (e.g. Zinger et al., 2018), often extends significantly beyond the dark matter splashback radius (Aung, Nagai & Lau, 2021). This results in a physically consequential region where the environment is already shaping galaxy properties. We therefore need a new tool that expands beyond a hard boundary and is instead sensitive to the physical properties of galaxies themselves. Our work introduces a probabilistic boundary designed precisely to mark this transitional zone using intrinsic galaxy properties.
@@ 2.2 Galaxy and cluster samples
The TNG simulation suite identifies groups using the Friends-of-Friends (FoF) algorithm (Davis et al., 1985), where particles are grouped together by their spatial distribution. The SubFind Algorithm (Springel et al., 2001; Dolag et al., 2009) identifies subhaloes based on gravitationally bound particles. The Friends-of-Friends (FoF) and SubFind algorithms are ran by the TNG Collaboration, which produce the group catalogue used in this work. From this catalogue, we first select samples out of all the groups and subhaloes at redshift $z=0$. The selection criteria is described as follows. Haloes with $M_{200,mean}$ (at $z=0$) > $10^{13}\,{\rm M}_{\odot}$ are considered clusters, and subhaloes with $M_{\star}$ > $10^{9}\,{\rm M}_{\odot}$ and $M_{dm}$ > $10^{9.5}\,{\rm M}_{\odot}$ are considered galaxies. The thresholds for the galaxy definitions was motivated by the resolution limit, using subhalo mass functions, as described in Nagai & Kravtsov (2005). We then group the cluster samples into four mass bins based on their $M_{200\rm,mean}$ at $z=0$: $10^{13\text{--}13.5}{\rm M}_{\odot}$, $10^{13.5\text{--}14}{\rm M}_{\odot}$, $10^{14\text{--}14.5}{\rm M}_{\odot}$, and $10^{14.5\text{--}15}{\rm M}_{\odot}$. These mass ranges are used for further analysis throughout the paper. They are chosen due to the lack of haloes with mass greater than $10^{15}{\rm M}_{\odot}$ to form a mass bin of $10^{15-15.5}{\rm M}_{\odot}$ . The mass bins were chosen to contain enough cluster samples for stacking, which prevented us from including haloes with mass greater than $10^{15}{\rm M}_{\odot}$.
Even though the simulation associates all subhaloes with host groups, these host groups may not suffice to be a cluster environment. Hence, we use a sorting algorithm to match each galaxy to the nearest cluster. The algorithm is applied to all subhaloes and matches them to our selected clusters. This way, we can ensure that the selected galaxies are paired with their true closest cluster within the box, instead of only relative to the set of chosen galaxies. We calculate the point-to-point distance in a periodic box. Algorithm 1 in Appendix A takes in cluster and subhalo positions in 3 dimensions and calculates the distance-to-closest cluster for each subhalo. The calculated distance is one dimensional, which is then normalised using the $R_{200,\rm mean}$ of the host cluster.